OpenStudy (kearsten):

solve using elimination. show steps plz 3x-7y+10=0, y-2x-3=0

OpenStudy (jdoe0001):

\( { \begin{array}{llll} 3x-7y+10=0\\ y-2x-3=0 \end{array}\implies \begin{array}{llll} 3x-7y+10=0\\ -2x+y-3=0 \end{array}\impliedby \begin{array}{llll} \textit{first, we arrange the "x" and}\\ \textit{"y" to be in the same column} \end{array} \\ \quad \\ \quad \\ \begin{array}{ccccll} 3x-7y+10=0&\implies &3x\cancel{-7y}+10=0\\ -2x+y-3=0&{\color{brown}{ \times 7}}\implies &-14x\cancel{+7y}-21=0 \\\hline\\ &&\square ?+\square ?=\square ? \end{array} }\) notice, arrange the "x" and "y" first so they're all in the same column, vertically

OpenStudy (jdoe0001):

we multiplied the bottom one, by 7 that way we end up with a +7y at the bottom that cancels out the -7y atop sum them up, vertically see what you end up with

OpenStudy (loser66):

\(\textit{wow, this guy uses another latex symbol} \implies \textit{there are many way to show math notations}\\\textit {wonder why he uses {1111} after each of array and why it doesn't work here} \\\textit{when I imitate his latex to get a gap between the context}\\t\textit {Which key did he use???}\)